Two dice are thrown. Find the probability that the sum of the 2 dice is greater than 7 given that the first die shows a 3.
Answer by stanbon(75887)

Two dice are thrown. Find the probability that the sum of the 2 dice is greater than 7 given that the first die shows a 3.
---
Winners: 3/5;3/6
---
Possible results with 1st being 3: 3/1;3/2;3/3;3/4;3/5;3/6
----
Probability(sum> 7|1st is 3) = 2/6 = 1/3
=============================================
Cheers,
Stan H.